OpenStudy (anonymous):

What is the slope of the line y = 3?

jimthompson5910 (jim_thompson5910):

Hint: y = 3 is the same as y = 0x + 3

jimthompson5910 (jim_thompson5910):

compare this to y = mx+b

myininaya (myininaya):

What kinda line is y=3? Is slope rise over run? Does the graph of y=3 have a rise? Does the graph of y=3 have a run?

OpenStudy (anonymous):

What is the slope of the line y = 3? undefined 0 –3 3

jimthompson5910 (jim_thompson5910):

another hint: y = mx+b has a slope of m

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

what does m equal in this case

jimthompson5910 (jim_thompson5910):

y = 0x+3 y = mx+b m = ??

OpenStudy (anonymous):

I got it now thank you :)

jimthompson5910 (jim_thompson5910):

what did you get